When point charges q1 = +5.7 μC and q2 = +2.8 μC are brought near each other, each experiences a repulsive force of magnitude 0.37 N. Determine the distance between the charges.
Answer #1

Force between two charges is given by:

F = \frac{1}{4\pi\epsilon_o}[\frac{qq'}{r^2}]

so, 0.37 = \frac{1}{4\pi\epsilon_o}[\frac{(5.7\times 10^{-6})(2.8\times 10^{-6})}{r^2}]

=> k = \frac{1}{4\pi\epsilon_o} = 9\times 10^9 Nm^2/C^2

therefore, 0.37 = 9\times 10^{9}[\frac{(5.7\times 10^{-6})(2.8\times 10^{-6})}{r^2}]

=> r = 0.623 m.